OpenStudy (anonymous):

What are the real or imaginary solutions of the polynomial equation? 27x^3+125=0 I have no idea how to do this

OpenStudy (radar):

It is the sum of two cubes:\[(3x)^{3} + 5^{3}\]

OpenStudy (anonymous):

What does that mean? I need help figuring out how to do this please

OpenStudy (radar):

It means that it is factorable and may be solved\[a ^{3}+b ^{3}= (a + b)(a ^{2}-ab + b ^{2}\]

OpenStudy (radar):

a=3x, b=5 (3x + 5)(9x^2 - 15x + 25)=0 thus 3x+5 = 0, x=-5/3 would be one answer, there will be 2 more solutions when you solve the quadratic.

OpenStudy (radar):

The other 2 solution are complex values, use the quadratic formula to solve the quadratic